OpenStudy (anonymous):

Really starting to dislike this site.. Asked this three times and can't understand it.. salesperson purchased a car 27 mi/gal city 38 mi/gal highway. Recent trip 1654 miles. Required 47 gallons. How many miles driven in city

sammixboo (sammixboo):

Here I'll help you better understand your problem :)

OpenStudy (anonymous):

thanks :)

OpenStudy (mathmale):

Is this the problem you mentioned in your message to me?

OpenStudy (anonymous):

yeah i have several others but this is one

sammixboo (sammixboo):

What is the required 47 gallons part?

OpenStudy (anonymous):

how many gallons it took for the trip

OpenStudy (mathmale):

Brit: I'll be right with you.

OpenStudy (anonymous):

i got to go get ready for work leave me a message and il check it later

OpenStudy (mathmale):

All right. I've solved the problem. You message me, please, when you're available to discuss this problem further.

OpenStudy (mathmale):

Looks like several helpers are offering you suggestions! I'll leave it to you, Brit, whether or not to contact me again to discuss this problem. Have a good day!

OpenStudy (texaschic101):

c + h = 47 --->h = 47 - c 27c + 38h = 1654 now sub 47 - c in for h in the 2nd equation 27c + 38(47 - c) = 1654 27c + 1786 - 38c = 1654 - 11c = 1654 - 1786 -11c = - 132 c = 132/11 c = 12 now sub 12 in for c c + h = 47 12 + h = 47 h = 47 - 12 h = 35 check... 27c + 38h = 1654 27(12) + 38(35) = 1654 324 + 1330 = 1654 1654 = 1654 (correct) 12 city miles were driven

OpenStudy (whpalmer4):

Drove 1654 miles using 47 gallons of gasoline. Highway miles are 38 to the gallon, city miles are 27 to the gallon. You can think of this problem in terms of how many gallons of gas are used in each sort of driving, and then find the miles from that at the end. Let's let \(x\) be the number of gallons used for city driving. Then \(47-x\) is the number of gallons used for highway driving. 27x = number of miles driven in the city 38(47-x) = number of miles driven on highway 27x+38(47-x) = 1654 total miles driven = city + highway \[27x + 38*47-38x = 1654\]Solve that for \(x\), and multiply \(x\) by 27 to get city miles.

OpenStudy (texaschic101):

crap....I messed up I found gallons instead of miles. So, whpalmer, would you multiply 12 x 27 to get the city miles ?

OpenStudy (whpalmer4):

Exactly!

OpenStudy (texaschic101):

I kinda went about finding the answer the long way

OpenStudy (mathmale):

TC: Appreciate your checking your own work and your honesty in admitting that your results were incorrect. Bet you can solve this problem!

OpenStudy (whpalmer4):

12*27 = 324 city miles (47-12)*38 = 1330 highway miles total = 1330+324 = 1654

OpenStudy (texaschic101):

you guys are so smart....glad you were here to find my mistake. THX

OpenStudy (mathmale):

@whpalmer4: Congrats. Your presentation is clear and accurate. My own work confirms the correctness of yours.

OpenStudy (whpalmer4):

You can write the equations to do it directly in miles, but then you end up having to do fractions: If we let \(c\) be city miles, then we use \(\frac{c}{27}\) gallons of gas on the city miles. Similarly, if \(h\) is highway miles, we use \(\frac{h}{38}\) gallons on the highway. \[\frac{c}{27} + \frac{h}{38} = 47\](total gallons used) \[c+h=1654\]Now solve by substitution, and get respective mileage quantities directly.